OpenStudy (anonymous):

3x^2+20/ 3x^2 + 4x - 5

OpenStudy (gorv):

\[3*x^{2}+\frac{ 20 }{ 3x^2 }+4x-5\]

OpenStudy (gorv):

is it ur equation

OpenStudy (anonymous):

Its for Algebra hw

OpenStudy (gorv):

no no wat i wrote that is ur Q????

OpenStudy (anonymous):

no. my equation is (3x^2+20)/ (3x^2 + 4x - 5). The entire expression (3x^2+20) is divided by (3x^2 + 4x - 5)

OpenStudy (gorv):

ok ok

OpenStudy (gorv):

\[\frac{ 3x^{2} +20+4x-4x}{ 3x^{2}+4x-5 }\]

OpenStudy (anonymous):

Oh ok, so then the terms cancel out and you divide 20 by -5 and your left with -4 -4x right?

OpenStudy (gorv):

\[\frac{ 3x^{2}+4x-5 -4x+20+5 }{ 3x^{2}+4x-5 }\]

OpenStudy (gorv):

\[\frac{ 3x^{2}+4x-5 }{ 3x^2 +4x-5}+\frac{ 25-4x }{ 3x^2+4x-5 }\]